### v1

# Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions

Abstract
This paper presents new constructions of sum-rank metric codes derived from algebraic function fields, while the existing results about such codes remain limited. A central challenge in this field involves the determination of parameters. We address this challenge through quadratic Kummer extensions, and propose two general constructions of $2\times 2$ sum-rank codes. The novel sum-rank metric code constructions derived from algebraic function fields demonstrate superior code lengths compared to conventional coding paradigms, particularly when contrasted with linearized Reed-Solomon codes under equivalent code parameter constraints. We also establish explicit parameters including dimensions and minimum distances of our codes. Finally, an illustrative example through elliptic function fields is constructed for validating the theoretical framework.
Index terms: Sum-rank metric codes, algebraic geometry codes, function fields, elliptic curves.
I Introduction
Algebraic function fields over finite fields provide a foundational framework for Hamming-metric code construction [15]. Unified methodological frameworks are constructed through exploitation of algebraic structures inherent in function fields, enabling simultaneous characterization of both minimum distance parameters and dual code invariants [41]. Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
$d_{H}:=\min_{c_{1}\neq c_{2}}\{wt_{H}(c_{1}-c_{2})\mid c_{1},c_{2}\in C\},$
where $wt_{H}$ counts non-zero coordinate positions. Classical algebraic geometry (AG) codes are constructed by evaluating functions in Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
For a linear $[n,k,d_{H}]$ code $C$, the Singleton bound asserts that
$d_{H}\leq n-k+1.$
When the equality holds, this code is called a maximum distance separable (MDS) code. Reed-Solomon (RS) codes [18], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent Hamming metric MDS codes. Their maximal length of $q+1$ over $\mathbb{F}_{q}$ coincides with the main conjecture of MDS codes.
Very recently, sum-rank metric codes have attracted significant attention due to applications in multishot network coding [22, 24, 31, 35], space-time coding [40], and distributed storage [23, 25]. For more applications and details, we refer the reader to [27]. The basic definitions are given as follows.
Definition I.1.
Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, denote
$\Pi:=\Pi_{q}(m_{1}\times n_{1}|\cdots|m_{s}\times n_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{m_{i}\times n_{i}}$
where $\mathbb{F}_{q}^{m_{i}\times n_{i}}$ denotes the space of $m_{i}\times n_{i}$ matrices over $\mathbb{F}_{q}$. Note that $\Pi$ forms an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}m_{i}n_{i}$. The sum-rank weight of
$X:=(X_{1},\ldots,X_{s})\in\Pi$
is defined as
$wt_{sr}(X):=\sum\limits_{i=1}^{s}rank(X_{i}).$
The sum-rank distance between $X,Y\in\Pi$ is
$d_{sr}(X,Y)=wt_{sr}(X-Y).$
The sum-rank weight and distance induce a metric on $\Pi$. We formally define a linear sum-rank metric code as follows.
Definition I.2.
A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$ with minimum sum-rank distance
$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in\Pi\}.$
For notational conciseness, we adopt the convention of referring to sum-rank metric codes as sum-rank codes throughout this work. The sum-rank code framework constitutes a natural generalization of classical coding systems: (i) When configured with block uniformity parameters $n_{i}=m_{i}=1$ for all $1\leq i\leq s$, these codes precisely recover the Hamming metric codes of length $s$ through coordinate-wise symbol comparison; (ii) In the monolithic regime where $s=1$, the construction collapses to the original rank-metric codes introduced by Delsarte [12]. This multifunctional parametrization reveals the intrinsic capacity of the codes to interpolate between symbol-level and matrix-level error protection strategies.
The theoretical landscape of sum-rank codes is underpinned by three foundational pillars as established in prior research:
  1. 1.
    Structural Properties: Code duality mechanisms and weight enumerator analysis [1, 5]
  2. 2.
    Distance Bounds: Fundamental limitations including Singleton-type and Johnson-type bounds [37]
  3. 3.
    Decoding Thresholds: Characterization of list-decoding capacity regions [2]
The Singleton bound for sum-rank codes was provided in [5] and [25]. For the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $10\leq\delta\leq n_{j}-$ 1. Then
$|C|\leq q^{\sum_{i=j}^{s}m_{i}n_{i}-m_{j}\delta}.$
A code $C\subset\Pi$ attaining this bound is called maximum sum-rank distance (MSRD). For the special case $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound is of the form
$|C|\leq q^{m(sn-d_{sr}+1)}.$
Throughout this paper, we focus on sum-rank metric codes with these uniform parameters.
Constructions of “good” sum-rank metric codes can be found in [1, 6, 7, 9, 23, 26, 28, 29, 30, 33, 34]. A standard technique constructs $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block. This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric. Moreover, efficient algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 16, 39].
One of the most prominent rank metric codes is the Gabidulin code, constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 14, 19, 21, 32]. Among sum-rank codes, linearized Reed-Solomon codes [23] represent the most extensively investigated families to date, see [8, 17, 20, 33, 38]. Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as rank metric and sum-rank metric analogues of RS codes.
Motivations
While existing algebraic geometric constructions [9, 28] for sum-rank metric codes are non-trivial, most results are constructed using towers of function fields as illustrated in Figure 1. In [4], E. Berardini and X. Caruso established the first geometric construction of sum-rank metric codes. Their approach employs the Ore polynomial ring [36] over an extension field derived from the completion of $F$ at specified places and subsequently develops the theory of Riemann-Roch spaces over this ring. This framework enables them to introduce linearized algebraic geometry codes and presents explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
A principal motivation for constructing codes via algebraic function fields lies in their ability to provide numerous evaluation places bounded by the Hasse-Weil bound with higher genus. Consequently, codes with long lengths can be constructed over small finite fields. Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [43]. Recent work [37] demonstrates that random linear sum-rank metric codes attain the asymptotic GV-like bound with high probability. This raises a natural question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
[…]
Figure 1: Tower of function field extensions.
Our contributions
In contrast to the construction in [4], we propose an alternative method to construct sum-rank metric codes via function fields. The key tool of our construction is the local integral basis of a place in function field extensions. Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and $F^{\prime}/F$ a finite separable function field extension. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ forms a vector space over $\mathcal{O}_{P}$. Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$. When $F^{\prime}/F$ is Galois, we establish a linearized-polynomial-like map on $\mathcal{O}^{\prime}_{P}$ through the direct sum of Riemann-Roch spaces, yielding sum-rank metric codes via evaluation of the function matrix at rational places. While our construction overlaps with [4], it differs substantially by emphasizing the intrinsic structure of function field extensions. By sacrificing certain subspaces in the direct sum of Riemann-Roch spaces, we explicitly determine the parameters of our sum-rank codes.
Organization
The remainder of this paper is organized as follows. Section II introduces the foundational concepts necessary for our purpose. In Section III, we establish a general construction of sum-rank metric codes through function field extensions. Section IV presents the explicit sum-rank metric codes via quadratic Kummer extensions and introduces a new specialized construction, including a case of elliptic function fields. Section V concludes this paper.
II Preliminaries and basic concepts
This section reviews some basic definitions and notations, primarily following [41].
II-A Places and divisors
Let $q$ be a prime power. An algebraic function field $F/\mathbb{F}_{q}$ denotes a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$. A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$ with
$\deg(P)=[\mathcal{O}_{P}/P:\mathbb{F}_{q}].$
If $\deg(P)=1$, then $P$ is called rational. The prime element of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{p}$. Any non-zero $z\in F$ has a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$. The associated discrete valuation
$v_{P}:F\to\mathbb{Z}\cup\{\infty\}$
satisfies $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality of $v_{P}$ is characterized as follows:
Lemma II.1.
For any discrete valuation $v_{P}$,
$v_{P}(x+y)\geq\min\{v_{P}(x),v_{P}(y)\}$
holds for all $x,y\in F$. Moreover, if $v_{P}(x)\neq v_{P}(y)$, then $v_{P}(x+y)=\min\{v_{P}(x),v_{P}(y)\}$.
A divisor $G$ is defined as the formal sum of places
$G:=\sum\limits_{P}n_{P}P$
with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}$. For any function $z\in F$,
$(z):=\sum\limits_{P}v_{P}(z)P$
is called the principal divisor. Given two divisors $G_{1},G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places. The associated Riemann-Roch space is defined as
$\mathcal{L}(G):=\{z\in F\mid(z)+G\geq 0\}\cup\{0\}.$
The space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
Lemma II.2.
[41, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
$\ell(G)\geq\deg(G)-g(F)+1.$
Moreover, if $\deg(G)\geq 2g(F)-1$, then
$\ell(G)=\deg(G)-g(F)+1.$
II-B Extensions of function fields
A finite algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $[F^{\prime}:F]<\infty$ and $m\geq 1$. A place $P^{\prime}$ of $F^{\prime}$ is said lying over $P$ if $P\subseteq P^{\prime}$, denoted $P^{\prime}|P$. The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
$v_{P^{\prime}}(x)=e(P^{\prime}|P)\cdot v_{P}(x)$
for all $x\in F$, while the relative degree of $P^{\prime}$ over $P$ is given by
$f(P^{\prime}|P)=[\mathcal{O}_{P^{\prime}}/P^{\prime}:\mathcal{O}_{P}/P].$
For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ with $P^{\prime}|P$. The conorm of $P$ is defined as
$\text{Con}_{F^{\prime}/F}(P):=\sum\limits_{P^{\prime}|P}e(P^{\prime}|P)\cdot P^{\prime}$
and
$\text{Con}_{F^{\prime}/F}\left(\sum\limits_{P}n_{P}P\right):=\sum n_{P}\cdot\text{Con}_{F^{\prime}/F}(P).$
Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ has the form
$(z)_{F^{\prime}}=\text{Con}_{F^{\prime}/F}((z)_{F}).$
The fundamental equality of places is stated as follows.
Proposition II.1.
[41, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$. Then
$\sum\limits_{i=1}^{n}e_{i}f_{i}=[F^{\prime}:F].$
If $[F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$. In this case, it follows that $e_{i}=f_{i}=1$. When $e(P^{\prime}|P)>1$ for some $P^{\prime}$, the place $P$ is ramified in $F^{\prime}/F$. Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
II-C Galois Extensions and Kummer Extensions
An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Let $\text{Gal}(F^{\prime}/F)$ be the Galois group of $F^{\prime}/F$. For any place $P$ of $F$, the Galois group acts transitively on the set of places lying over $P$.
Kummer extensions constitute a special class of Galois extensions. We present their definition and key properties.
Definition II.1.
Let $\mathbb{F}_{q}$ contain a primitive $n$-th root of unity $\zeta$. Suppose that $u\in F$ satisfies
$u\neq\omega^{d}$
for all $\omega\in F$ and $1<d|n$. Let $F^{\prime}=F(y)$ with $y^{n}=u$ be an extension of $F/\mathbb{F}_{q}$. Then $F^{\prime}$ is called a Kummer extension of $F$ with:
  • $\text{Gal}(F^{\prime}/F)=\langle\sigma\rangle$ where $\sigma(y)=\zeta y$;
  • For any place $P$ of $F$ and $P^{\prime}$ lying over $P$, $e(P^{\prime}|P)=\frac{n}{\gcd(n,v_{P}(u))}$;
  • If there exists a place $Q$ of $F$ with $\gcd(n,v_{Q}(u))=1$, then $\mathbb{F}_{q}$ remains the constant field of $F^{\prime}$, and
    $g(F^{\prime})=1+n(g+1)+\frac{1}{2}\sum\limits_{P}(n-\gcd(n,v_{P}(u)))\deg(P);$
  • For any place $P$ with $v_{P}(u))\geq 0$, the basis $\{1,y,\ldots,y^{n-1}\}$ is a local integral basis of $F^{\prime}/F$ for $P$.
III A General Setting of Sum-Rank Metric Codes
In this section, we present a general construction of sum-rank metric codes through function field extensions, which differs from the approach by E. Berardini and X. Caruso in [4].
III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
Let $F/\mathbb{F}_{q}$ be a function field and $F^{\prime}/\mathbb{F}_{q^{m}}$ a Galois extension of $F/\mathbb{F}_{q}$ with $[F^{\prime}:F]=n$. Suppose that $P$ is a place of $F$ and $t\in F$ is a prime element of $P$. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
$\mathcal{O}^{\prime}_{P}=\bigcap\limits_{i=1}^{n}\mathcal{O}_{P_{i}}.$
An integral basis $\{z_{1},\ldots,z_{n}\}$ of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is a basis of $F^{\prime}/F$ satisfying
$\mathcal{O}^{\prime}_{P}=\sum\limits_{i=1}^{n}\mathcal{O}_{P}z_{i}.$
The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over $\mathcal{O}_{P}/P$ with basis
$\{z_{1}+t\mathcal{O}^{\prime}_{P},\ldots,z_{n}+t\mathcal{O}^{\prime}_{P}\}.$
For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and place $P^{\prime}$ with $P^{\prime}|P$, observe that for $z\in\mathcal{O}^{\prime}_{P}$:
$v_{P^{\prime}}(\sigma(z))=v_{\sigma^{-1}(P^{\prime})}(z)\geq 0$
since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, hence $\sigma(V)=V$. The set
$\{\sigma(z_{1})+t\mathcal{O}^{\prime}_{P},\ldots,\sigma(z_{n})+t\mathcal{O}^{\prime}_{P}\}$
forms another basis of $V$ over $\mathcal{O}_{P}/P$. Then $\sigma$ induces a matrix $\bar{A}_{\sigma}$ over $\mathcal{O}_{P}/P$ with respect to $\{z_{i}\}$.
Similarly, for $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$. This induces an $\mathcal{O}_{P}/P$-linear map $\bar{\mu}_{f}:V\to V$ via
$\bar{\mu}_{f}(z+t\mathcal{O}^{\prime}_{P}):=fz+t\mathcal{O}^{\prime}_{P}.$
This induces a matrix $\bar{B}_{f}$ with entries in $\mathcal{O}_{P}/P$.
Remark III.1.
The matrix $A=(a_{ij})$ associated to $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
$\pi:\mathcal{O}_{P}\to\mathcal{O}_{P}/P,$
we obtain $\bar{A}=(\pi(a_{ij}))$. This explains the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
When $P$ is rational (i.e., $\mathcal{O}_{P}/P=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}l(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$. Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$. We define a linear operator
$L:\mathcal{O}^{\prime}_{P}$ $\to\mathcal{O}^{\prime}_{P}$
$g$ $\mapsto\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g).$
This $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ satisfying
$\varepsilon_{P}(L)=\sum\limits_{i=1}^{n}A_{i}B_{i}=(M_{ij})$
where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Furthermore, let
$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij})).$
We establish the following lemma.
Lemma III.1.
  • (i)
    Let $\{z_{1},\ldots,z_{n}\}$ be an integral basis of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$. Then the matrix representing $L$ with respect to this basis is $\varepsilon_{P}(L)$.
  • (ii)
    The determinant $\det(\varepsilon_{P}(L))$ belongs to $\mathcal{O}_{P}$.
  • (iii)
    The kernel ${\rm ker}(L)\neq\{0\}$ if and only if $\det(\varepsilon_{P}(L))=0$.
Proof.
(i) Let $g=g_{1}z_{1}+\cdots+g_{n}z_{n}\in\mathcal{O}^{\prime}_{P}$. For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}_{P}$, let $A_{\sigma}$ and $B_{f}$ denote the matrices defined above. Then
$\mu_{f}(g)=[g_{1},\ldots,g_{n}]B_{f}[z_{1},\ldots,z_{n}]^{T}$
and
$\sigma(g)=[g_{1},\ldots,g_{n}]A_{\sigma}[z_{1},\ldots,z_{n}]^{T}.$
Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The conclusion follows from the linearity of $\mu_{f}$ and $\sigma$.
(ii) The result immediately follows since all coefficients of $A_{\sigma_{i}}$ and $B_{f_{i}}$ are contained in $\mathcal{O}_{P}$.
(iii) Consider the equation
$\varepsilon_{P}(L)^{T}[x_{1},\ldots,x_{n}]^{T}=[0,\ldots,0].$
Any solution $[g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$. Therefore ${\rm ker}(L)$ contains non-zero elements if and only if the system has non-trivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$. ∎
III-B Sum-rank Codes in General Case
Now we present the construction of sum-rank metric codes. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Assume that the divisor $Q_{0}$ is totally ramified in $F^{\prime}/F$ with $P_{0}|Q_{0}$. Let $G=nQ_{0}$, $G^{\prime}=nP_{0}$ and $D=Q_{1}+\cdots+Q_{s}$. Define the evaluation map:
$\text{ev}_{D,1}:\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$. Thus $\varepsilon_{Q_{i}}(L)$ are well-defined. Moreover, the function space $\mathcal{L}(G^{\prime})$ forms an $\mathbb{F}_{q}$-linear space and $L$ is $F$-linear. It follows that $\text{ev}_{D,1}$ is an $\mathbb{F}_{q}$-linear map. Thus the image of $\text{ev}_{D,1}$ is a subspace of $\Pi$. The code $C_{sr,1}(D,G)$ is defined as the image of $\text{ev}_{D,1}$.
Remark III.2.
The code $C_{sr,1}$ resembles the code $C(1,G,Q_{1},\ldots,Q_{s})$ defined in [4, Definition 4]. In their work, E. Berardini and X. Caruso constructed codes through the isomorphism chain
$D_{F^{\prime}_{i},x}\to D_{F^{\prime}_{i},1}\to\text{End}_{F_{i}}(F^{\prime}_{i})$
where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, while $F^{\prime}_{i}$ represents the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,n$. Our approach directly employs $L$ to express linearized-polynomial-like maps, corresponding to elements of $D_{F^{\prime},1}$. We subsequently utilize the integral closure $\mathcal{O}^{\prime}_{Q_{i}}$ to construct explicit matrices $\bar{\varepsilon}_{Q_{i}}(L)$, where $\varepsilon_{Q_{i}}(L)$ is induced by $L$ at $Q_{i}$.
Furthermore, while previous work proposed several hypotheses for code parameter determination, the ring $D_{F^{\prime}_{i},1}$ fails to satisfy all necessary conditions. In subsequent sections, we provide more explicit dimension and minimum distance bounds by considering subspaces of $\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$.
Proposition III.1.
The length $n^{2}s$ of $C_{sr,1}$ over $\mathbb{F}_{q}$ satisfies
$n^{2}s\leq n^{2}(q+2g(F)\sqrt{q}).$
Proof.
In the construction of $C_{sr,1}$, each rational place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$. By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ contains at most $q+1+2g(F)\sqrt{q}$ rational places, implying $s\leq q+2g(F)\sqrt{q}$. ∎
IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
IV-A Parameters of the General Construction
We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance for $C_{sr,1}$.
Lemma IV.1.
Suppose that $Q\neq P$ is a rational place of $F$ with $Q$ totally ramified in $F^{\prime}/F$. Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$. For $f_{1},\ldots,f_{n}\in\mathcal{L}(kQ^{\prime})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists $f_{i}$ satisfying
$v_{Q^{\prime}}(f_{i})<v_{Q^{\prime}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ are contained in $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$. The discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ therefore exceeds the minimal valuation in its determinant expansion. Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$. Since $Q$ is totally ramified, the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=nQ^{\prime}.$
By Lemma III.1 (ii) and the conorm properties, we have
$v_{Q}(\det(\varepsilon_{P}(L)))\leq k.$
Note that for any $Q^{\prime}$ of $F^{\prime}$, there exists only one $Q$ in $F$ with $Q^{\prime}|Q$, thus $\det(\varepsilon_{P}(L))$ only has poles at $Q$. Therefore $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$. For any non-zero $g\in\mathcal{O}^{\prime}_{P}$, the equality
$v_{Q^{\prime}}(\sigma_{i}(g))=v_{Q^{\prime}}(\sigma_{1}(g))$
holds for $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ when $2\leq j\leq n$. This yields
$v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(f_{j}\sigma_{j}(g)).$
By the strict triangle inequality, we have
$v_{Q^{\prime}}\left(\sum\limits_{j=1}^{n}f_{j}\sigma_{j}(g)\right)=v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(\sigma_{1}(g)).$
Thus $L(g)\neq 0$ for $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
Lemma IV.2.
Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated tp $Q\neq P$ while $rank(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
Proof.
The condition $rank(\bar{\varepsilon}_{P}(L))<n$ means $\det(\bar{\varepsilon}_{P_{\alpha}}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
$\bar{\varepsilon}_{P}(L)=(\pi_{P}(M_{ij}))=(M_{ij})(P).$
Therefore
$\det(\bar{\varepsilon}_{P}(L))=\det(\varepsilon_{P}(L))(P).$
Since $\det(\varepsilon_{P}(L))(P)=0$, it follows that $\det(\varepsilon_{P}(L))\in P$. ∎
Lemma IV.3.
Let $k,k_{1}$ be integers with $k>k_{1}\geq 2g(F)-1$, and $P$ a rational place of $F$. Define
$\mathcal{L}^{*}_{k_{1}}(kP):=\{f|-k\leq v_{P}(f)<-k_{1}\}\cup\{0\}.$
Then $\mathcal{L}^{*}_{k_{1}}(kP)$ forms a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$.
Proof.
By the Riemann-Roch theorem, we obtain that $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions $\ell_{k}:=k-g(F)+1$ and $\ell_{k_{1}}:=k_{1}-g(F)+1$, respectively. Thus,
$\mathcal{L}^{*}_{k_{1}}(kP):=\mathcal{L}(kP)\setminus\mathcal{L}(k_{1}P)\cup\{0\}.$
For $g_{1},g_{2}\in\mathcal{L}^{*}_{k_{1}}(kP)$ and $a\in\mathbb{F}_{q}$, we observe that $g_{1}+g_{2}\in\mathcal{L}(kP)$ and $ag_{1}\in\mathcal{L}^{*}_{k_{1}}(kP)$. It remains to show $g_{1}+g_{2}\notin\mathcal{L}(k_{1}P)$.
Let $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$ be a basis for $\mathcal{L}(kP)$. Since $k_{1}\geq 2g-1$, the space $\mathcal{L}(k_{1}P)$ has a basis contained in $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$. Suppose that
$g_{1}+g_{2}=\sum\limits_{i=\ell_{k_{1}}+1}^{\ell_{k}}a_{i}\omega_{i}$
with $a_{i}\in\mathbb{F}_{q}$. It follows that $g_{1}+g_{2}\in\mathcal{L}(k_{1}P)$ implies $g_{1}+g_{2}=a_{1}\omega_{1}+\cdots+a_{\ell_{k_{1}}}\omega_{\ell_{k_{1}}}$. The linear independence of basis elements forces all coefficients to vanish, which means $g_{1}+g_{2}=0\in\mathcal{L}^{*}_{k_{1}}(kP)$. ∎
Theorem IV.1.
Suppose that $F^{\prime}/F$ is a Kummer extension of degree 2 with a totally ramified place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$. Let $s,k,k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0}),f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})\}$
has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
Proof.
The code $C_{sr,1}$ has length exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
$k-k_{1}+\ell_{k_{1}}-\dim(\mathcal{L}(kQ_{\infty}-\text{Con}_{F^{\prime}/F}(\sum\limits_{i=1}^{s}Q_{s}))).$
Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, which implies that the dimension of $C_{sr,1}$ simplifies to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$. Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$. Then the codeword of $L$ has sum-rank distance $d_{L}=\sum\limits_{i=1}^{s}d_{i}$. By Lemma IV.2, we have
$v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
for each $i$. Furthermore, Lemma IV.1 shows
$\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, we get $\det(\varepsilon_{P_{\alpha}}(L))\neq 0$. When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &f_{22}u,&f_{21}\end{bmatrix}$ (1)
where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then It follows from $\det(\varepsilon_{Q_{i}}(L))=0$ that $u-f_{*}^{2}$ belong to $F$ for a certain function $f_{*}$. Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish as $u\neq\omega^{2}$ for any $\omega\in F$. Therefore $\det(\varepsilon_{P_{\alpha}}(L))$ has at most $k$ zeros. This implies that
$\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore we have $d_{sr}\geq 2s-k$. This completes the whole proof. ∎
IV-B New Construction for Quadratic Extension
For a quadratic extension $F^{\prime}/F$, we provide an alternative construction by utilizing completely split places. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Consider a place $Q_{0}$ that splits completely in $F^{\prime}/F$ with $Q_{0,1}$ and $Q_{0,2}$ lying over $Q_{0}$. Let $G=nQ,G_{i}=nQ_{0,i}$ and $D=Q_{1}+\cdots+Q_{s}$. The evaluation map is constructed as follows:
$\text{ev}_{D,2}:\bigoplus_{i=1}^{2}\mathcal{L}(G_{i})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
The image of $\text{ev}_{D,2}$ forms a linear subspace of $\Pi$. Subsequently, the code $C_{sr,2}(D,G)$ is defined as the image. To analyze the parameters of $C_{sr,2}(D,G)$, we require the following lemma.
Lemma IV.4.
Let $Q\neq P$ be a rational divisor of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists $f_{i}$ such that
$v_{Q_{i}}(f_{i})<v_{Q_{j}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
First, every term in the determinant expansion of $\varepsilon_{P}(L)$ must be contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=Q_{1}+Q_{2}.$
Following a similar argument to Lemma IV.1, we obtain
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Second, w.l.o.g, assume that $v_{Q_{1}}(f_{1})<v_{Q_{2}}(f_{2})$. For any non-zero function $g$, we have
$v_{Q_{2}}(\sigma_{2}(g))=v_{Q_{1}}(\sigma_{1}(g)).$
If $L(g)=0$, then
$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))<v_{Q_{1}}(\sigma_{1}(g)).$
This implies
$v_{Q_{2}}(\sigma_{1}(g))<v_{Q_{2}}(\sigma_{2}(g))$
and consequently
$v_{Q_{2}}(f_{1}\sigma_{1}(g))<v_{Q_{2}}(f_{2}\sigma_{2}(g)).$
It follows that $v_{Q_{2}}(L(g))<\infty$, contradicting $L(g)=0$. Hence $L(g)\neq 0$ for all non-zero $g\in\mathcal{O}^{\prime}_{P}$. The conclusion follows from Lemma III.1 (iii). ∎
The subsequent result parallels Theorem IV.1, thus we omit the details of the proof.
Theorem IV.2.
Suppose that $F^{\prime}/F$ is a function field extension of degree 2. Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1}|Q_{0}$ and $Q_{0,2}|Q_{0}$. Let $s,k,k_{1}$ be integers such that $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
IV-C Examples from Elliptic Function Fields
An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ of genus 1. This serves as a canonical example since evaluation places can be selected from all $q+1$ places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$, limiting the code length to $q$. Nevertheless, this suffices to illustrate our framework.
In this section, we consider the case when $q$ is odd, as elliptic function fields over even characteristic are Artin-Schreier extensions. Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
$y^{2}=f_{\mathcal{E}}$
with $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ being a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
$\sigma(y)=-y.$
Let $P_{\infty}$ be the place at infinity of $\mathbb{F}_{q}(x)$, which is totally ramified in $\mathcal{E}/\mathbb{F}_{q}(x)$. Suppose that $Q_{\infty}$ is the place of $\mathcal{E}$ lying over $P_{\infty}$. The basis $\{1,y\}$ of $\mathcal{E}/\mathbb{F}_{q}(x)$ forms an integral basis for all places of $\mathcal{E}/\mathbb{F}_{q}$ except $P_{\infty}$. Consequently, for any rational place $P\neq P_{\infty}$, the matrix induced by $\sigma$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}1,&0\\ 0,&-1\end{bmatrix}.$
Let $f_{1}$ and $f_{2}$ be two functions over $\mathcal{E}/\mathbb{F}_{q}$, where each $f_{i}$ has the form $f_{i}=f_{i1}+f_{i2}y$ with $f_{i1}$ and $f_{i2}$ belonging to $\mathbb{F}_{q}(x)$ for $i=1,2$. For any rational place $P\neq P_{\infty}$, the matrix induced by $f_{i}$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}f_{i1},&f_{i2}\\ f_{i2}f_{\mathcal{E}},&f_{i1}\end{bmatrix}.$
Let $L=f_{1}+f_{2}\sigma$. Then we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}f_{11}+f_{21},&f_{12}+f_{22}\\ (f_{12}-f_{22})f_{\mathcal{E}},&f_{11}-f_{21}\end{bmatrix}$
and
$\det(\varepsilon_{P}(L))=f_{11}^{2}-f_{21}^{2}-(f_{12}^{2}-f_{22}^{2})f_{\mathcal{E}}.$
For positive integers $k,k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{\infty}),f_{2}\in\mathcal{L}(k_{1}Q_{\infty})\}$
and
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}\in\mathbb{F}_{q}(x)/\mathbb{F}_{q}$. Through the Singleton bound combined with Theorems IV.1 and IV.2, we derive
$4s-2k\leq 2d_{sr,1},2d_{sr,2}\leq 4s-k+2,$
where $s\leq q$. The following two examples are given by SageMath [42].
Example IV.1.
Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
$\{P_{0},P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$
where $P_{\alpha}$ corresponds to the prime element $x-\alpha$. For $k=6$ and $k_{1}=3$, the space $\mathcal{L}(6Q_{\infty})$ has basis $\{1,x,y,x^{2},xy,x^{3}\}$. The code $C_{sr,1}(D,6Q_{\infty})$ has parameters $[24,6,6]$. Taking $f_{1}=3$ and $f_{2}=y$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}3,&1\\ -(x^{3}+3),&3\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=x^{3}+5.$
Since $x^{3}+5$ is irreducible over $\mathbb{F}_{q}[x]$, the function $\det(\varepsilon_{P}(L))$ has no rational zeros. Consequently, the codeword
$\left(\begin{bmatrix}3,&1\\ -3,&3\end{bmatrix},\ldots,\begin{bmatrix}3,&1\\ 5,&3\end{bmatrix}\right)$
achieves the maximum sum-rank distance $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}x^{3}+1,&0\\ 0,&1-x^{3}\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=1-x^{6}.$
All places in $\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$ are zeros of $\det(\varepsilon_{P}(L))$. Thus, the codeword
$\left(\begin{bmatrix}1,&0\\ 0,&1\end{bmatrix},\ldots,\begin{bmatrix}0,&0\\ 0,&2\end{bmatrix}\right)$
has sum-rank distance 8, attaining the lower bound.
Example IV.2.
Consider the same finite field $\mathbb{F}_{q}$ and elliptic curve $\mathcal{E}$ as in the previous example. We choose the place $P_{1}$, and let $Q_{1,1}$ and $Q_{1,2}$ be the places of $\mathcal{E}/\mathbb{F}_{q}$ lying over $P_{1}$. We take the rational places
$\{P_{0},P_{2},P_{3},P_{4},P_{5},P_{6}\}.$
Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
$A_{0}=1,A_{6}=36,A_{7}=144,A_{8}=1542,A_{9}=7944,$
$A_{10}=26904,A_{11}=46959,A_{12}=34122$
where $A_{i}$ counts the number of codewords with sum-rank weight $i$.
V Conclusion
In this paper, we conducted two families of $2\times 2$ sum-rank metric codes by exploiting quadratic Kummer extensions of algebraic function fields. By representing elements in the extension fields as matrices over the base field through explicit algebraic mappings, we rigorously determined the parameters of the proposed codes, including their minimum sum-rank distances and code rates. To demonstrate the practical applicability of our framework, we provided concrete instantiations over elliptic function fields.
To the best of our knowledge, these code constructions are novel in the context of algebraic-geometric (AG) codes for the sum-rank metric. Prior to this work, the only systematic study of AG codes in this metric was established by Berardini and Caruso [4], who proposed a general framework for such codes. However, our methodology differs fundamentally from theirs by explicitly exploiting the intrinsic algebraic structures of function field extensions—particularly through matrix representations derived from quadratic Kummer extensions—to control code parameters and optimize distance properties.
In this paper, we investigate Kummer extensions defined by the equation $y^{2}=u\quad\text{where }u\notin\{\omega^{2}\mid\omega\in F\setminus\mathbb{F}_{q}\}.$ This specific choice of non-square elements $u$ enables us to rigorously establish the non-degeneracy condition $\det(\varepsilon_{P}(L))\neq 0$ in the restricted case where $f_{1}=0$. The principal challenge in generalizing these constructions to $n\times n$ matrices stems from the need to preserve the non-vanishing determinant property $\det(\varepsilon_{P}(L))\neq 0\quad\text{for arbitrary }n\geq 2.$ We intend to address this dimensionality constraint through more flexible algebraic frameworks in subsequent work.
Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension $k-g(F)+1$, we observe the following approximate relationship:
$\max d_{sr}-\min d_{sr}\approx\frac{k+g(F)+1}{2}.$
This raises an interesting problem: whether techniques developed for constructing MDS AG codes [11] can be adapted to the sum-rank metric setting to achieve MSRD AG codes. Following the approach for AG codes with narrow bounds (such as elliptic codes [11]), we propose this problem as an important objective for our subsequent research.
VI Acknowledgment
The work of Chang-An Zhao is supported by Guangdong Basic and Applied Basic Research Foundation(No. 2025A1515011764), Guangdong Major Project of Basic and Applied Basic Research (No. 2019B030302008), the National Natural Science Foundation of China (No. 12441107) and Guangdong Provincial Key Laboratory of Information Security Technology(No. 2023B1212060026).

### v2

@@
-Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions
+Constructing Sum-Rank Metric Codes from Quadratic Galois Extensions of Function Fields
 Abstract
-This paper presents new constructions of sum-rank metric codes derived from algebraic function fields, while the existing results about such codes remain limited.
-A central challenge in this field involves the determination of parameters.
-We address this challenge through quadratic Kummer extensions, and propose two general constructions of $2\times 2$ sum-rank codes.
-The novel sum-rank metric code constructions derived from algebraic function fields demonstrate superior code lengths compared to conventional coding paradigms, particularly when contrasted with linearized Reed-Solomon codes under equivalent code parameter constraints.
-We also establish explicit parameters including dimensions and minimum distances of our codes.
-Finally, an illustrative example through elliptic function fields is constructed for validating the theoretical framework.
+This paper introduces new constructions of sum-rank metric codes derived from algebraic function fields, as existing results on such codes remain limited.
+A major challenge lies in the determination of their parameters.
+We address this issue by employing quadratic Galois extensions, proposing two general constructions of $2\times 2$ sum-rank codes.
+Analogous to algebraic geometry codes in the Hamming metric, our codes achieve a larger block length compared to existing constructions.
+We determine explicit parameters including dimensions and minimum distances of our codes, and we present an illustrative example using elliptic function fields.
+Finally, we discuss the asymptotic behavior of our codes and compare them with the Gilbert-Varshamov-like bound for sum-rank metric codes.
 Index terms: Sum-rank metric codes, algebraic geometry codes, function fields, elliptic curves.
 I Introduction
-Algebraic function fields over finite fields provide a foundational framework for Hamming-metric code construction [15].
-Unified methodological frameworks are constructed through exploitation of algebraic structures inherent in function fields, enabling simultaneous characterization of both minimum distance parameters and dual code invariants [41].
+Algebraic function fields over finite fields provide a fundamental framework for modern error-correcting codes in the Hamming metric [16], as the function field perspective enables a deeper analysis of code parameters, including minimum distance and duality properties [43].
 Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
 $d_{H}:=\min_{c_{1}\neq c_{2}}\{wt_{H}(c_{1}-c_{2})\mid c_{1},c_{2}\in C\},$
-where $wt_{H}$ counts non-zero coordinate positions.
-Classical algebraic geometry (AG) codes are constructed by evaluating functions in Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
-For a linear $[n,k,d_{H}]$ code $C$, the Singleton bound asserts that
+where $wt_{H}$ denotes the Hamming weight, counting non-zero coordinate positions.
+Classical algebraic geometry (AG) codes are constructed by evaluating functions from Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
+For a linear $n,k,d_{H}]$ code $C$, the Singleton bound states that
 $d_{H}\leq n-k+1.$
-When the equality holds, this code is called a maximum distance separable (MDS) code.
-Reed-Solomon (RS) codes [18], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent Hamming metric MDS codes.
-Their maximal length of $q+1$ over $\mathbb{F}_{q}$ coincides with the main conjecture of MDS codes.
-Very recently, sum-rank metric codes have attracted significant attention due to applications in multishot network coding [22, 24, 31, 35], space-time coding [40], and distributed storage [23, 25].
-For more applications and details, we refer the reader to [27].
-The basic definitions are given as follows.
+When equality holds, the code is called a maximum distance separable (MDS) code.
+Reed-Solomon (RS) codes [19], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent examples of Hamming metric MDS codes.
+Their maximum length of $q+1$ over $\mathbb{F}_{q}$ aligns with the main conjecture of MDS codes.
+Recently, sum-rank metric codes have garnered significant attention due to their applications in multishot network coding [24, 26, 33, 37], space-time coding [42], and distributed storage [25, 27].
+For further details and applications, we refer the reader to [29].
+The basic definitions are provided as follows.
 Definition I.1.
-Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, denote
-$\Pi:=\Pi_{q}(m_{1}\times n_{1}|\cdots|m_{s}\times n_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{m_{i}\times n_{i}}$
-where $\mathbb{F}_{q}^{m_{i}\times n_{i}}$ denotes the space of $m_{i}\times n_{i}$ matrices over $\mathbb{F}_{q}$.
-Note that $\Pi$ forms an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}m_{i}n_{i}$.
+Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, define
+$\Pi:=\Pi_{q}(n_{1}\times m_{1}|\cdots|n_{s}\times m_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n_{i}\times m_{i}}$
+where $\mathbb{F}_{q}^{n_{i}\times m_{i}}$ denotes the space of $n_{i}\times m_{i}$ matrices over $\mathbb{F}_{q}$.
+Note that $\Pi$ is an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}n_{i}m_{i}$.
 The sum-rank weight of
@@
 is defined as
-$wt_{sr}(X):=\sum\limits_{i=1}^{s}rank(X_{i}).$
-The sum-rank distance between $X,Y\in\Pi$ is
+$wt_{sr}(X):=\sum\limits_{i=1}^{s}{\rm rank}(X_{i}).$
+The sum-rank distance between $X$ and $Y$ in $\Pi$ is
 $d_{sr}(X,Y)=wt_{sr}(X-Y).$
-The sum-rank weight and distance induce a metric on $\Pi$. We formally define a linear sum-rank metric code as follows.
+The sum-rank weight and distance induce a metric on $\Pi$. We present the definition as follows.
 Definition I.2.
-A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$ with minimum sum-rank distance
-$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in\Pi\}.$
-For notational conciseness, we adopt the convention of referring to sum-rank metric codes as sum-rank codes throughout this work.
-The sum-rank code framework constitutes a natural generalization of classical coding systems: (i) When configured with block uniformity parameters $n_{i}=m_{i}=1$ for all $1\leq i\leq s$, these codes precisely recover the Hamming metric codes of length $s$ through coordinate-wise symbol comparison; (ii) In the monolithic regime where $s=1$, the construction collapses to the original rank-metric codes introduced by Delsarte [12].
-This multifunctional parametrization reveals the intrinsic capacity of the codes to interpolate between symbol-level and matrix-level error protection strategies.
-The theoretical landscape of sum-rank codes is underpinned by three foundational pillars as established in prior research:
-  1. 1.
-    Structural Properties: Code duality mechanisms and weight enumerator analysis [1, 5]
-  2. 2.
-    Distance Bounds: Fundamental limitations including Singleton-type and Johnson-type bounds [37]
-  3. 3.
-    Decoding Thresholds: Characterization of list-decoding capacity regions [2]
-The Singleton bound for sum-rank codes was provided in [5] and [25].
-For the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $10\leq\delta\leq n_{j}-$ 1.
+A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$. The minimum sum-rank distance of $C_{sr}$ is defined as
+$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in C_{sr}\}.$
+Sum-rank metric codes generalize Hamming metric codes of length $s$ when $n_{i}=m_{i}=1$ for all $1\leq i\leq s$.
+When $s=1$, they reduce to rank metric codes, which were first introduced by Delsarte [12].
+Fundamental properties and bounds for sum-rank metric codes are examined in [1, 2, 5, 39].
+The Singleton bound was established in [5] and [27].
+In the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $0\leq\delta\leq n_{j}-1$.
 Then
 $|C|\leq q^{\sum_{i=j}^{s}m_{i}n_{i}-m_{j}\delta}.$
-A code $C\subset\Pi$ attaining this bound is called maximum sum-rank distance (MSRD). For the special case $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound is of the form
+A code $C\subseteq\Pi$ that attains this bound is called a maximum sum-rank distance (MSRD) code.
+For the special case where $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound takes the form
 $|C|\leq q^{m(sn-d_{sr}+1)}.$
-Throughout this paper, we focus on sum-rank metric codes with these uniform parameters.
-Constructions of “good” sum-rank metric codes can be found in [1, 6, 7, 9, 23, 26, 28, 29, 30, 33, 34].
-A standard technique constructs $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block.
+Throughout this paper, we concentrate on sum-rank metric codes with these uniform parameters.
+Let $N:=sn$. For fixed positive real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
+$\mathcal{R}_{sr}<$ $\delta_{sr}^{2}-\delta_{sr}(2+\frac{2}{N})+1+\frac{2}{N}+\frac{1}{N^{2}}$
+$-\frac{\sum_{i=1}^{d_{sr}N-1}\log_{q}(1+\frac{s-1}{i})+\log_{q}(\delta_{sr}N-1)}{Nm}-\frac{\log_{q}(\gamma_{q})}{nm},$
+where $\gamma_{q}=\prod_{i=1}^{\infty}(1-q^{-i})^{-1}$, there exists a sum-rank code with rate at least $\mathcal{R}_{sr}$ and relative minimum sum-rank distance at least $\delta_{sr}$.
+When $m=\xi n$ goes to the infinity and $m\in\omega(\log_{q}(s))$, where $\xi$ is a constant, we have
+$\mathcal{R}_{sr}\sim\delta_{sr}^{2}-\delta_{sr}(1+\frac{1}{\xi})+1.$
+Recent work [39] demonstrates that random linear sum-rank metric codes attain the asymptotic Gilbert-Varshamov-like (GV-like) bound with high probability.
+Consider the case where $n=m$ and the number of blocks $s$ goes to infinity.
+Then the previous bound is asymptotically provided by [4]:
+$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{n}\log_{q}(1+\frac{1}{\delta_{sr}n})-\frac{\log_{q}(1+\delta_{sr}n)}{n^{2}}$
+$-\frac{\log_{q}(\gamma_{q})}{n^{2}}+o(1).$
+Constructions of “good” sum-rank metric codes are available in several works [1, 6, 7, 9, 25, 28, 30, 31, 32, 35, 36].
+A standard technique involves constructing $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block.
 This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric.
-Moreover, efficient algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 16, 39].
-One of the most prominent rank metric codes is the Gabidulin code, constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 14, 19, 21, 32].
-Among sum-rank codes, linearized Reed-Solomon codes [23] represent the most extensively investigated families to date, see [8, 17, 20, 33, 38].
-Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as rank metric and sum-rank metric analogues of RS codes.
+Moreover, algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 17, 41].
+One of the most prominent rank metric codes is the Gabidulin code, which is constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 15, 21, 23, 34].
+Among sum-rank metric codes, linearized Reed-Solomon codes [25] represent the most extensively investigated family to date, see [8, 18, 22, 35, 40].
+Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as generalizations of RS codes.
 Motivations
-While existing algebraic geometric constructions [9, 28] for sum-rank metric codes are non-trivial, most results are constructed using towers of function fields as illustrated in Figure 1.
+While non-trivial algebraic geometric constructions for sum-rank metric codes exist, see [9, 30], most results are constructed using towers of function fields as illustrated below.
 In [4], E.
@@
 Caruso established the first geometric construction of sum-rank metric codes.
-Their approach employs the Ore polynomial ring [36] over an extension field derived from the completion of $F$ at specified places and subsequently develops the theory of Riemann-Roch spaces over this ring.
-This framework enables them to introduce linearized algebraic geometry codes and presents explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
-A principal motivation for constructing codes via algebraic function fields lies in their ability to provide numerous evaluation places bounded by the Hasse-Weil bound with higher genus.
+Their approach employs the Ore polynomial ring [38] over an extension field derived from the completion of $F$ at specified places, and subsequently develops the theory of Riemann-Roch spaces over this ring.
+This framework allows them to define linearized algebraic geometry codes and present explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
+A principal motivation for constructing codes via algebraic function fields lies in the availability of numerous evaluation places bounded by the Hasse-Weil bound for higher genus.
 Consequently, codes with long lengths can be constructed over small finite fields.
 Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [45].
-Recent work [37] demonstrates that random linear sum-rank metric codes attain the asymptotic GV-like bound with high probability.
-This raises a natural question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
+This naturally raises the question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
 […]
@@
 Our contributions
-In contrast to the construction in [4], we propose an alternative method to construct sum-rank metric codes via function fields.
-The key tool of our construction is the local integral basis of a place in function field extensions.
-Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and $F^{\prime}/F$ a finite separable function field extension.
-The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ forms a vector space over $\mathcal{O}_{P}$.
+In contrast to the construction in [4], we propose an approach for constructing sum-rank metric codes using function fields.
+The central element of our construction is the local integral basis at a place in function field extensions.
+Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and let $F^{\prime}/F$ be a finite separable function field extension.
+The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ constitutes a vector space over $\mathcal{O}_{P}$.
 Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$.
-When $F^{\prime}/F$ is Galois, we establish a linearized-polynomial-like map on $\mathcal{O}^{\prime}_{P}$ through the direct sum of Riemann-Roch spaces, yielding sum-rank metric codes via evaluation of the function matrix at rational places.
-While our construction overlaps with [4], it differs substantially by emphasizing the intrinsic structure of function field extensions.
-By sacrificing certain subspaces in the direct sum of Riemann-Roch spaces, we explicitly determine the parameters of our sum-rank codes.
+When $F^{\prime}/F$ is Galois, we define a map on $\mathcal{O}^{\prime}_{P}$, via the direct sum of Riemann-Roch spaces.
+This yields sum-rank metric codes through the evaluation of the function matrix at rational places.
+Although our construction shares some similarities with [4], it significantly differs by focusing on the intrinsic structure of function field extensions.
+By decomposing the Riemann-Roch spaces as a direct sum of subspaces, we explicitly determine the parameters of our sum-rank codes.
+Furthermore, when $q$ is a special square, we analyze the asymptotic behavior of our codes and compare them with the GV-like bound.
 Organization
 The remainder of this paper is organized as follows.
-Section II introduces the foundational concepts necessary for our purpose.
-In Section III, we establish a general construction of sum-rank metric codes through function field extensions.
-Section IV presents the explicit sum-rank metric codes via quadratic Kummer extensions and introduces a new specialized construction, including a case of elliptic function fields.
+Section II introduces the fundamental concepts required for our purpose.
+In Section III, we develop a general construction of sum-rank metric codes via function field extensions.
+Section IV provides the explicit sum-rank metric codes using quadratic Galois extensions and presents a new specialized construction, including an instance involving elliptic function fields and the comparison with the GV-like bound.
 Section V concludes this paper.
 II Preliminaries and basic concepts
-This section reviews some basic definitions and notations, primarily following [41].
+This section reviews fundamental definitions and notations, primarily based on [43].
 II-A Places and divisors
 Let $q$ be a prime power.
-An algebraic function field $F/\mathbb{F}_{q}$ denotes a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$.
-A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$ with
-$\deg(P)=[\mathcal{O}_{P}/P:\mathbb{F}_{q}].$
+An algebraic function field $F/\mathbb{F}_{q}$ is a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$.
+A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$.
+Denote by $\hat{F}_{P}:=\mathcal{O}_{P}/P$ the residue field of $P$, which is a finite extension of $\mathbb{F}_{q}$.
+The degree of $P$ is defined as
+$\deg(P)=[\hat{F}_{P}:\mathbb{F}_{q}].$
 If $\deg(P)=1$, then $P$ is called rational.
-The prime element of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{p}$.
-Any non-zero $z\in F$ has a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$.
-The associated discrete valuation
+The prime element (or uniformizing parameter) of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{P}$.
+Any non-zero element $z\in F$ admits a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$.
+The discrete valuation associated with $P$ is
 $v_{P}:F\to\mathbb{Z}\cup\{\infty\}$
-satisfies $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality of $v_{P}$ is characterized as follows:
+, satisfying $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality is characterized as follows:
 Lemma II.1.
-For any discrete valuation $v_{P}$,
+For any discrete valuation $v_{P}$ with associated place $P$,
 $v_{P}(x+y)\geq\min\{v_{P}(x),v_{P}(y)\}$
 holds for all $x,y\in F$. Moreover, if $v_{P}(x)\neq v_{P}(y)$, then $v_{P}(x+y)=\min\{v_{P}(x),v_{P}(y)\}$.
-A divisor $G$ is defined as the formal sum of places
+A divisor $G$ is defined as a formal sum of places
 $G:=\sum\limits_{P}n_{P}P$
-with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}$. For any function $z\in F$,
+with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}\deg(P)$. The support of $G$ is defined as
+$\text{supp}(G):=\{P\mid n_{P}\neq 0\}.$
+For any function $z\in F$,
 $(z):=\sum\limits_{P}v_{P}(z)P$
-is called the principal divisor. Given two divisors $G_{1},G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places. The associated Riemann-Roch space is defined as
+is called the principal divisor of $z$.
+Given two divisors $G_{1}$ and $G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places $P$.
+The Riemann-Roch space associated with $G$ is defined as
 $\mathcal{L}(G):=\{z\in F\mid(z)+G\geq 0\}\cup\{0\}.$
-The space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
+This space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
 Lemma II.2.
-[41, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
+[43, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
 $\ell(G)\geq\deg(G)-g(F)+1.$
@@
 II-B Extensions of function fields
-A finite algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $[F^{\prime}:F]<\infty$ and $m\geq 1$.
-A place $P^{\prime}$ of $F^{\prime}$ is said lying over $P$ if $P\subseteq P^{\prime}$, denoted $P^{\prime}|P$.
+An algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $F^{\prime}:F]<\infty$ and $m\geq 1$.
+A place $P^{\prime}$ of $F^{\prime}$ is said lying over a place $P$ of $F$ if $P\subseteq P^{\prime}$, denoted by $P^{\prime}|P$.
 The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
 $v_{P^{\prime}}(x)=e(P^{\prime}|P)\cdot v_{P}(x)$
-for all $x\in F$, while the relative degree of $P^{\prime}$ over $P$ is given by
+for all $x\in F$, and the relative degree of $P^{\prime}$ over $P$ is given by
 $f(P^{\prime}|P)=[\mathcal{O}_{P^{\prime}}/P^{\prime}:\mathcal{O}_{P}/P].$
-For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ with $P^{\prime}|P$. The conorm of $P$ is defined as
+For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ such that $P^{\prime}|P$. The conorm of $P$ is defined as
 $\text{Con}_{F^{\prime}/F}(P):=\sum\limits_{P^{\prime}|P}e(P^{\prime}|P)\cdot P^{\prime}$
@@
 $\text{Con}_{F^{\prime}/F}\left(\sum\limits_{P}n_{P}P\right):=\sum n_{P}\cdot\text{Con}_{F^{\prime}/F}(P).$
-Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ has the form
+Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ satisfies
 $(z)_{F^{\prime}}=\text{Con}_{F^{\prime}/F}((z)_{F}).$
@@
 Proposition II.1.
-[41, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$. Then
+[43, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$ for $i=1,\ldots,n$. Then
 $\sum\limits_{i=1}^{n}e_{i}f_{i}=[F^{\prime}:F].$
-If $[F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$.
-In this case, it follows that $e_{i}=f_{i}=1$.
-When $e(P^{\prime}|P)>1$ for some $P^{\prime}$, the place $P$ is ramified in $F^{\prime}/F$.
+If $F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$.
+In this case, $e_{i}=f_{i}=1$ for all $i$.
+If $e(P^{\prime}|P)>1$ for some $P^{\prime}$, then $P$ is ramified in $F^{\prime}/F$.
 Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
-II-C Galois Extensions and Kummer Extensions
-An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension.
-Let $\text{Gal}(F^{\prime}/F)$ be the Galois group of $F^{\prime}/F$.
-For any place $P$ of $F$, the Galois group acts transitively on the set of places lying over $P$.
-Kummer extensions constitute a special class of Galois extensions. We present their definition and key properties.
-Definition II.1.
-Let $\mathbb{F}_{q}$ contain a primitive $n$-th root of unity $\zeta$. Suppose that $u\in F$ satisfies
-$u\neq\omega^{d}$
-for all $\omega\in F$ and $1<d|n$. Let $F^{\prime}=F(y)$ with $y^{n}=u$ be an extension of $F/\mathbb{F}_{q}$. Then $F^{\prime}$ is called a Kummer extension of $F$ with:
-  • $\text{Gal}(F^{\prime}/F)=\langle\sigma\rangle$ where $\sigma(y)=\zeta y$;
-  • For any place $P$ of $F$ and $P^{\prime}$ lying over $P$, $e(P^{\prime}|P)=\frac{n}{\gcd(n,v_{P}(u))}$;
-  • If there exists a place $Q$ of $F$ with $\gcd(n,v_{Q}(u))=1$, then $\mathbb{F}_{q}$ remains the constant field of $F^{\prime}$, and
-    $g(F^{\prime})=1+n(g+1)+\frac{1}{2}\sum\limits_{P}(n-\gcd(n,v_{P}(u)))\deg(P);$
-  • For any place $P$ with $v_{P}(u))\geq 0$, the basis $\{1,y,\ldots,y^{n-1}\}$ is a local integral basis of $F^{\prime}/F$ for $P$.
+II-C Galois Extensions
+An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Specifically, the automorphism group
+[…]
+has order $F^{\prime}:F]$. This group is called the Galois group and write $\text{Gal}(F^{\prime}/F):=\text{Aut}(F^{\prime}/F)$.
+For any place $P$ of $F/\mathbb{F}_{q}$, the Galois group $\text{Gal}(F^{\prime}/F)$ acts transitively on the set of places
+[…]
+via $\sigma(P^{\prime})=\{\sigma(x)\mid x\in P^{\prime}\}$. For any place $P^{\prime}$ of $F^{\prime}/\mathbb{F}_{q^{m}}$, the valuation corresponding to the place $\sigma(P^{\prime})$ is given by
+$v_{\sigma(P^{\prime})}(y)=v_{P^{\prime}}(\sigma^{-1}(y)),\quad y\in F^{\prime}.$
 III A General Setting of Sum-Rank Metric Codes
-In this section, we present a general construction of sum-rank metric codes through function field extensions, which differs from the approach by E. Berardini and X. Caruso in [4].
+In this section, we introduce a general construction of sum-rank metric codes using function field extensions, which differs from the approach proposed by E. Berardini and X. Caruso in [4].
 III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
-Let $F/\mathbb{F}_{q}$ be a function field and $F^{\prime}/\mathbb{F}_{q^{m}}$ a Galois extension of $F/\mathbb{F}_{q}$ with $[F^{\prime}:F]=n$.
-Suppose that $P$ is a place of $F$ and $t\in F$ is a prime element of $P$.
+Let $F/\mathbb{F}_{q}$ be a function field, and $F^{\prime}/\mathbb{F}_{q^{m}}$ be a Galois extension of $F/\mathbb{F}_{q}$ with $F^{\prime}:F]=n$.
+Suppose that $P$ is a place of $F$, and $t\in F$ is a prime element of $P$.
 The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
@@
 $\mathcal{O}^{\prime}_{P}=\sum\limits_{i=1}^{n}\mathcal{O}_{P}z_{i}.$
-The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over $\mathcal{O}_{P}/P$ with basis
+The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over the finite field $\hat{F}_{P}$ with basis
 $\{z_{1}+t\mathcal{O}^{\prime}_{P},\ldots,z_{n}+t\mathcal{O}^{\prime}_{P}\}.$
-For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and place $P^{\prime}$ with $P^{\prime}|P$, observe that for $z\in\mathcal{O}^{\prime}_{P}$:
+For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and a place $P^{\prime}$ with $P^{\prime}|P$, observe that for any $z\in\mathcal{O}^{\prime}_{P}$,
 $v_{P^{\prime}}(\sigma(z))=v_{\sigma^{-1}(P^{\prime})}(z)\geq 0$
-since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, hence $\sigma(V)=V$. The set
+since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, and hence $\sigma(V)=V$. The set
 $\{\sigma(z_{1})+t\mathcal{O}^{\prime}_{P},\ldots,\sigma(z_{n})+t\mathcal{O}^{\prime}_{P}\}$
-forms another basis of $V$ over $\mathcal{O}_{P}/P$. Then $\sigma$ induces a matrix $\bar{A}_{\sigma}$ over $\mathcal{O}_{P}/P$ with respect to $\{z_{i}\}$.
-Similarly, for $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$.
-This induces an $\mathcal{O}_{P}/P$-linear map $\bar{\mu}_{f}:V\to V$ via
+forms another basis of $V$ over $\hat{F}_{P}$. Then $\sigma$ induces an $n\times n$ matrix $\bar{A}_{\sigma}$ over $\hat{F}_{P}$ with respect to the basis $\{z_{i}\}$.
+Similarly, for any $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication map $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$.
+This induces an $\hat{F}_{P}$-linear map $\bar{\mu}_{f}:V\to V$ via
 $\bar{\mu}_{f}(z+t\mathcal{O}^{\prime}_{P}):=fz+t\mathcal{O}^{\prime}_{P}.$
-This induces a matrix $\bar{B}_{f}$ with entries in $\mathcal{O}_{P}/P$.
+This also induces an $n\times n$ matrix $\bar{B}_{f}$ with entries in $\hat{F}_{P}$.
 Remark III.1.
-The matrix $A=(a_{ij})$ associated to $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
-$\pi:\mathcal{O}_{P}\to\mathcal{O}_{P}/P,$
-we obtain $\bar{A}=(\pi(a_{ij}))$. This explains the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
-When $P$ is rational (i.e., $\mathcal{O}_{P}/P=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}l(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$.
+The matrix $A=(a_{ij})$ associated with $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
+$\pi:\mathcal{O}_{P}\to\hat{F}_{P},$
+we obtain $\bar{A}=(\pi(a_{ij}))$. This justifies the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
+When $P$ is rational (i.e., $\hat{F}_{P}=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$.
 Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$.
@@
 $g$ $\mapsto\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g).$
-This $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ satisfying
+For any two functions $g_{1},g_{2}\in F$, we have
+$L(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}(g_{1}+g_{2})=L(g_{1})+L(g_{2}).$
+Thus, this $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ such that
 $\varepsilon_{P}(L)=\sum\limits_{i=1}^{n}A_{i}B_{i}=(M_{ij})$
-where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Furthermore, let
-$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij})).$
-We establish the following lemma.
+where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively.
+Then $\varepsilon_{P}(L)$ is an $n\times n$ matrix over $\mathcal{O}_{P}$.
+Moreover, let
+$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij}))$
+be the matrix over $\mathbb{F}_{q}$ induced by $\varepsilon_{P}(L)$. We now state the following lemma.
 Lemma III.1.
@@
 $\sigma(g)=[g_{1},\ldots,g_{n}]A_{\sigma}[z_{1},\ldots,z_{n}]^{T}.$
-Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The conclusion follows from the linearity of $\mu_{f}$ and $\sigma$.
+Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The result follows from the linearity of the maps $\mu_{f}$ and $\sigma$.
 (ii) The result immediately follows since all coefficients of $A_{\sigma_{i}}$ and $B_{f_{i}}$ are contained in $\mathcal{O}_{P}$.
@@
 $\varepsilon_{P}(L)^{T}[x_{1},\ldots,x_{n}]^{T}=[0,\ldots,0].$
-Any solution $[g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$.
-Therefore ${\rm ker}(L)$ contains non-zero elements if and only if the system has non-trivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$.
+Any solution $g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$.
+Therefore ${\rm ker}(L)$ contains nonzero elements if and only if the system has nontrivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$.
 ∎
-III-B Sum-rank Codes in General Case
-Now we present the construction of sum-rank metric codes.
-Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$.
-Assume that the divisor $Q_{0}$ is totally ramified in $F^{\prime}/F$ with $P_{0}|Q_{0}$.
-Let $G=nQ_{0}$, $G^{\prime}=nP_{0}$ and $D=Q_{1}+\cdots+Q_{s}$.
+III-B Sum-rank Codes in the General Case
+We now present the construction of sum-rank metric codes.
+Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$, and $Q_{0}$ be another place of $F$.
+Assume that $P_{0,1},\ldots,P_{0,m}$ are all places of $F^{\prime}$ lying over $Q_{0}$.Let $G=nQ_{0}$, $D=Q_{1}+\cdots+Q_{s}$, and $G_{j}$ are $n$ divisor of $F^{\prime}$ with
+$\text{supp}(G_{j})\subset\{P_{0,1},\ldots,P_{0,m}\}.$
 Define the evaluation map:
-$\text{ev}_{D,1}:\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$ $\to\Pi$
-$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
+$\text{ev}_{D}:\mathcal{L}(G_{1})\oplus\cdots\oplus\mathcal{L}(G_{n})$ $\to\Pi=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n\times n}$
+$(f_{1},\ldots,f_{n})$ $\mapsto(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))$
+where $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$.
 Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$.
 Thus $\varepsilon_{Q_{i}}(L)$ are well-defined.
-Moreover, the function space $\mathcal{L}(G^{\prime})$ forms an $\mathbb{F}_{q}$-linear space and $L$ is $F$-linear.
-It follows that $\text{ev}_{D,1}$ is an $\mathbb{F}_{q}$-linear map.
-Thus the image of $\text{ev}_{D,1}$ is a subspace of $\Pi$.
-The code $C_{sr,1}(D,G)$ is defined as the image of $\text{ev}_{D,1}$.
+Note that $\mathcal{L}(G_{j})$ is an $\mathbb{F}_{q}$-linear space.
+Since $L$ is $F$-linear, it follows that $\text{ev}_{D}$ is an $\mathbb{F}_{q}$-linear map.
+Thus the image of $\text{ev}_{D}$ is a subspace of $\Pi$.
+The code $C_{sr}(D,G)$ is defined as the image of $\text{ev}_{D}$, which is exactly a linear space over $\mathbb{F}_{q}$.
 Remark III.2.
-The code $C_{sr,1}$ resembles the code $C(1,G,Q_{1},\ldots,Q_{s})$ defined in [4, Definition 4]. In their work, E. Berardini and X. Caruso constructed codes through the isomorphism chain
+In the work of [4], E. Berardini and X. Caruso constructed codes through the isomorphism chain
 $D_{F^{\prime}_{i},x}\to D_{F^{\prime}_{i},1}\to\text{End}_{F_{i}}(F^{\prime}_{i})$
-where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, while $F^{\prime}_{i}$ represents the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,n$.
-Our approach directly employs $L$ to express linearized-polynomial-like maps, corresponding to elements of $D_{F^{\prime},1}$.
-We subsequently utilize the integral closure $\mathcal{O}^{\prime}_{Q_{i}}$ to construct explicit matrices $\bar{\varepsilon}_{Q_{i}}(L)$, where $\varepsilon_{Q_{i}}(L)$ is induced by $L$ at $Q_{i}$.
-Furthermore, while previous work proposed several hypotheses for code parameter determination, the ring $D_{F^{\prime}_{i},1}$ fails to satisfy all necessary conditions.
-In subsequent sections, we provide more explicit dimension and minimum distance bounds by considering subspaces of $\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$.
+where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, and $F^{\prime}_{i}$ denotes the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,s$.
+The functor for generating the matrix in [4] has form $\sum\limits_{i=0}^{n-1}f_{i}\Phi^{i}$, where $\Phi^{r}=x$ for some $0\neq x\in F$, and $f_{i}\in\Lambda_{L,x}(G^{\prime})$.
+To determine the code parameters, E.
+Berardini and X.
+Caruso hypothesized that $D_{L,x}$ has no nonzero zero divisor.
+Our approach directly employs the operator $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$, corresponding to elements of $D_{F^{\prime},1}$ since $F^{\prime}/F$ is Galois.
+However, the ring $D_{F^{\prime},1}$ is not a domain because the polynomial $T^{r}-1\in F^{\prime}[T]$ is reducible.
+For the determination of code parameters, we consider $Q_{0}$ that is totally ramified or splits completely in $F^{\prime}/F$, and then present another choice for $f_{i}$, which is explicitly described in subsequent sections.
 Proposition III.1.
 The length $n^{2}s$ of $C_{sr,1}$ over $\mathbb{F}_{q}$ satisfies
-$n^{2}s\leq n^{2}(q+2g(F)\sqrt{q}).$
+$n^{2}s\leq n^{2}(q+2g(F)\sqrt{q}+1).$
 Proof.
-In the construction of $C_{sr,1}$, each rational place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$.
-By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ contains at most $q+1+2g(F)\sqrt{q}$ rational places, implying $s\leq q+2g(F)\sqrt{q}$.
+In the construction of $C_{sr}$, each place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$.
+By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ has at most $q+1+2g(F)\sqrt{q}$ rational places, which implies that $s\leq q+2g(F)\sqrt{q}+1$.
 ∎
 IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
-IV-A Parameters of the General Construction
-We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance for $C_{sr,1}$.
+IV-A Construction with Totally Ramified Rational Places
+We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance.
 Lemma IV.1.
-Suppose that $Q\neq P$ is a rational place of $F$ with $Q$ totally ramified in $F^{\prime}/F$.
+Let $P$ be a rational place of $F$.
+Suppose that $Q\neq P$ is a place of $F$ such that $Q$ is totally ramified in $F^{\prime}/F$.
 Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$.
@@
 $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
-Furthermore, if there exists $f_{i}$ satisfying
+Furthermore, if there exists an $f_{i}$ satisfying
 $v_{Q^{\prime}}(f_{i})<v_{Q^{\prime}}(f_{j}),\forall j\neq i,$
@@
 Proof.
-Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ are contained in $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$.
-The discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ therefore exceeds the minimal valuation in its determinant expansion.
+Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ belong to $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$.
+Therefore, the discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ is greater than the minimal valuation in its determinant expansion.
 Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$.
-Since $Q$ is totally ramified, the conorm of $Q$ is
+As $Q$ is totally ramified, the conorm of $Q$ is
 $\text{Con}_{F^{\prime}/F}(Q)=nQ^{\prime}.$
-By Lemma III.1 (ii) and the conorm properties, we have
+By Lemma III.1 (ii) and the properties of the conorm, we obtain
 $v_{Q}(\det(\varepsilon_{P}(L)))\leq k.$
-Note that for any $Q^{\prime}$ of $F^{\prime}$, there exists only one $Q$ in $F$ with $Q^{\prime}|Q$, thus $\det(\varepsilon_{P}(L))$ only has poles at $Q$.
-Therefore $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$.
-For any non-zero $g\in\mathcal{O}^{\prime}_{P}$, the equality
+Note that for any place $Q^{\prime}$ of $F^{\prime}$, there is a unique place $Q$ in $F$ such that $Q^{\prime}|Q$, so $\det(\varepsilon_{P}(L))$ has poles only at $Q$.
+Hence, $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$.
+For any nonzero $g\in\mathcal{O}^{\prime}_{P}$, the equality
 $v_{Q^{\prime}}(\sigma_{i}(g))=v_{Q^{\prime}}(\sigma_{1}(g))$
-holds for $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ when $2\leq j\leq n$. This yields
+holds for all $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ for all $j=2,\ldots,n$. This implies that
 $v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(f_{j}\sigma_{j}(g)).$
-By the strict triangle inequality, we have
+By the strict triangle inequality, it follows that
 $v_{Q^{\prime}}\left(\sum\limits_{j=1}^{n}f_{j}\sigma_{j}(g)\right)=v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(\sigma_{1}(g)).$
-Thus $L(g)\neq 0$ for $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
+Thus $L(g)\neq 0$ for any $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
 Lemma IV.2.
-Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated tp $Q\neq P$ while $rank(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
+Let $P$ be a place of $F$ with a prime element $t$.
+If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated to $Q\neq P$ while ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
 Proof.
-The condition $rank(\bar{\varepsilon}_{P}(L))<n$ means $\det(\bar{\varepsilon}_{P_{\alpha}}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
+The condition ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$ means that $\det(\bar{\varepsilon}_{P}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
 $\bar{\varepsilon}_{P}(L)=(\pi_{P}(M_{ij}))=(M_{ij})(P).$
@@
 Since $\det(\varepsilon_{P}(L))(P)=0$, it follows that $\det(\varepsilon_{P}(L))\in P$. ∎
-Lemma IV.3.
-Let $k,k_{1}$ be integers with $k>k_{1}\geq 2g(F)-1$, and $P$ a rational place of $F$. Define
-$\mathcal{L}^{*}_{k_{1}}(kP):=\{f|-k\leq v_{P}(f)<-k_{1}\}\cup\{0\}.$
-Then $\mathcal{L}^{*}_{k_{1}}(kP)$ forms a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$.
+Let $k,k_{1}$ be integers such that $k>k_{1}\geq 2g(F)-1$, and $P$ be a rational place of $F$.
+Then by the Riemann-Roch theorem, we obtain $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions
+$\ell_{k}:=k-g(F)+1,\ell_{k_{1}}:=k_{1}-g(F)+1,$
+respectively. Let
+$\mathcal{L}(kP)=\langle\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\rangle$
+where $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k_{1}}}\}$ forms a basis for $\mathcal{L}(k_{1}P)$. We define
+$\mathcal{L}^{*}_{k_{1}}(kP):=\langle\omega_{\ell_{k_{1}}+1},\ldots,\omega_{\ell_{k}}\rangle.$
+It is obvious that $\mathcal{L}^{*}_{k_{1}}(kP)$ is a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$. Then we have:
+Theorem IV.1.
+Suppose that $F^{\prime}/F$ is a Galois extension of degree 2 with a totally ramified rational place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$.
+Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$.
+For distinct places $Q_{1},\ldots,Q_{s}$, the code
+$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0}),f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})\}$
+has parameters $4s,\ell_{k},d_{sr}\geq 2s-k]$.
 Proof.
-By the Riemann-Roch theorem, we obtain that $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions $\ell_{k}:=k-g(F)+1$ and $\ell_{k_{1}}:=k_{1}-g(F)+1$, respectively.
-Thus,
-$\mathcal{L}^{*}_{k_{1}}(kP):=\mathcal{L}(kP)\setminus\mathcal{L}(k_{1}P)\cup\{0\}.$
-For $g_{1},g_{2}\in\mathcal{L}^{*}_{k_{1}}(kP)$ and $a\in\mathbb{F}_{q}$, we observe that $g_{1}+g_{2}\in\mathcal{L}(kP)$ and $ag_{1}\in\mathcal{L}^{*}_{k_{1}}(kP)$.
-It remains to show $g_{1}+g_{2}\notin\mathcal{L}(k_{1}P)$.
-Let $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$ be a basis for $\mathcal{L}(kP)$.
-Since $k_{1}\geq 2g-1$, the space $\mathcal{L}(k_{1}P)$ has a basis contained in $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$.
-Suppose that
-$g_{1}+g_{2}=\sum\limits_{i=\ell_{k_{1}}+1}^{\ell_{k}}a_{i}\omega_{i}$
-with $a_{i}\in\mathbb{F}_{q}$.
-It follows that $g_{1}+g_{2}\in\mathcal{L}(k_{1}P)$ implies $g_{1}+g_{2}=a_{1}\omega_{1}+\cdots+a_{\ell_{k_{1}}}\omega_{\ell_{k_{1}}}$.
-The linear independence of basis elements forces all coefficients to vanish, which means $g_{1}+g_{2}=0\in\mathcal{L}^{*}_{k_{1}}(kP)$.
-∎
-Theorem IV.1.
-Suppose that $F^{\prime}/F$ is a Kummer extension of degree 2 with a totally ramified place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$.
-Let $s,k,k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$.
-For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
-$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0}),f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})\}$
-has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
-Proof.
-The code $C_{sr,1}$ has length exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
+The length of $C_{sr,1}$ is exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
 $k-k_{1}+\ell_{k_{1}}-\dim(\mathcal{L}(kQ_{\infty}-\text{Con}_{F^{\prime}/F}(\sum\limits_{i=1}^{s}Q_{s}))).$
-Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, which implies that the dimension of $C_{sr,1}$ simplifies to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
+Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, implying that the dimension of $C_{sr,1}$ reduces to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
 Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$.
 Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$.
-Then the codeword of $L$ has sum-rank distance $d_{L}=\sum\limits_{i=1}^{s}d_{i}$.
+Then the sum-rank weight of the codeword corresponding to $L$ is $d_{L}=\sum\limits_{i=1}^{s}d_{i}$.
 By Lemma IV.2, we have
 $v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
-for each $i$. Furthermore, Lemma IV.1 shows
+for each $i$. Furthermore, Lemma IV.1 shows that
 $\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
-From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, we get $\det(\varepsilon_{P_{\alpha}}(L))\neq 0$. When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
-$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &f_{22}u,&f_{21}\end{bmatrix}$ (1)
-where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$.
-Then It follows from $\det(\varepsilon_{Q_{i}}(L))=0$ that $u-f_{*}^{2}$ belong to $F$ for a certain function $f_{*}$.
-Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish as $u\neq\omega^{2}$ for any $\omega\in F$.
-Therefore $\det(\varepsilon_{P_{\alpha}}(L))$ has at most $k$ zeros.
+From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, it follows that $\det(\varepsilon_{Q_{i}}(L))\neq 0$.
+Denote $F^{\prime}=F(y)$.
+Since $F^{\prime}:F]=2$, there exists a monic irreducible quadratic polynomial $\phi(T)\in F[T]$ that is the minimal polynomial of $y$ over $F$, write $\phi(T)=T^{2}+a_{1}T+a_{2}$.
+When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
+$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &-f_{22}a_{2},&f_{21}-f_{22}a_{1}\end{bmatrix}$
+where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then $\det(\varepsilon_{Q_{i}}(L))=0$ means
+$f_{21}^{2}-f_{21}f_{22}a_{1}+f_{22}^{2}a_{2}=0.$
+If $f_{22}=0$, the previous equation implies $f_{21}=0$ and $f_{2}=0$.
+If $f_{21}\neq 0$, then the function $\frac{-f_{21}}{f_{22}}$ is a root of $\phi(T)$.
+Both cases lead to a contradiction.
+Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish, and it has at most $k$ zeros.
 This implies that
 $\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
-Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore we have $d_{sr}\geq 2s-k$. This completes the whole proof. ∎
-IV-B New Construction for Quadratic Extension
-For a quadratic extension $F^{\prime}/F$, we provide an alternative construction by utilizing completely split places.
-Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$.
-Consider a place $Q_{0}$ that splits completely in $F^{\prime}/F$ with $Q_{0,1}$ and $Q_{0,2}$ lying over $Q_{0}$.
-Let $G=nQ,G_{i}=nQ_{0,i}$ and $D=Q_{1}+\cdots+Q_{s}$.
-The evaluation map is constructed as follows:
-$\text{ev}_{D,2}:\bigoplus_{i=1}^{2}\mathcal{L}(G_{i})$ $\to\Pi$
-$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
-The image of $\text{ev}_{D,2}$ forms a linear subspace of $\Pi$.
-Subsequently, the code $C_{sr,2}(D,G)$ is defined as the image.
-To analyze the parameters of $C_{sr,2}(D,G)$, we require the following lemma.
-Lemma IV.4.
-Let $Q\neq P$ be a rational divisor of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
+Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore, we have $d_{sr}\geq 2s-k$. ∎
+IV-B Construction with Splitting Completely Rational Places
+For a Galois extension $F^{\prime}/F$ of degree 2, the following lemmas are required to determine code parameters.
+Lemma IV.3.
+Let $P$ be a rational place of $F$.
+Suppose that $Q\neq P$ is a place of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ are the places of $F^{\prime}$ lying over $Q$.
+For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
 $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
-Furthermore, if there exists $f_{i}$ such that
+Furthermore, if there exists an $f_{i}$ satisfying
 $v_{Q_{i}}(f_{i})<v_{Q_{j}}(f_{j}),\forall j\neq i,$
@@
 Proof.
-First, every term in the determinant expansion of $\varepsilon_{P}(L)$ must be contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
+First, every term in the determinant expansion of $\varepsilon_{P}(L)$ is contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
 $\text{Con}_{F^{\prime}/F}(Q)=Q_{1}+Q_{2}.$
-Following a similar argument to Lemma IV.1, we obtain
+By an argument similar to that in Lemma IV.1, we obtain
 $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
@@
 $v_{Q_{2}}(\sigma_{2}(g))=v_{Q_{1}}(\sigma_{1}(g)).$
-If $L(g)=0$, then
-$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))<v_{Q_{1}}(\sigma_{1}(g)).$
+If $L(g)=0$, that is, $f_{1}\sigma_{1}(g)+f_{2}\sigma_{2}(g)=0$. Then by the strict triangle inequality and $v_{Q_{1}}(f_{1})\leq 0$, it follows that
+$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))\leq v_{Q_{1}}(\sigma_{1}(g)).$
 This implies
-$v_{Q_{2}}(\sigma_{1}(g))<v_{Q_{2}}(\sigma_{2}(g))$
+$v_{Q_{2}}(\sigma_{1}(g))\leq v_{Q_{2}}(\sigma_{2}(g))$
 and consequently
@@
 It follows that $v_{Q_{2}}(L(g))<\infty$, contradicting $L(g)=0$. Hence $L(g)\neq 0$ for all non-zero $g\in\mathcal{O}^{\prime}_{P}$. The conclusion follows from Lemma III.1 (iii). ∎
-The subsequent result parallels Theorem IV.1, thus we omit the details of the proof.
+The subsequent result is analogous to Theorem IV.1, so we omit proof details.
 Theorem IV.2.
 Suppose that $F^{\prime}/F$ is a function field extension of degree 2.
-Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1}|Q_{0}$ and $Q_{0,2}|Q_{0}$.
-Let $s,k,k_{1}$ be integers such that $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$.
-For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
+Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1},Q_{0,2}$ lying over $Q_{0}$.
+Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$.
+For distinct places $Q_{1},\ldots,Q_{s}$, the code
 $C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
-has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
+has parameters $4s,\ell_{k},d_{sr}\geq 2s-k]$.
 IV-C Examples from Elliptic Function Fields
-An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ of genus 1.
-This serves as a canonical example since evaluation places can be selected from all $q+1$ places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$, limiting the code length to $q$.
-Nevertheless, this suffices to illustrate our framework.
-In this section, we consider the case when $q$ is odd, as elliptic function fields over even characteristic are Artin-Schreier extensions.
+An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ and has genus 1.
+It serves as a canonical example and suffices to illustrate our framework.
+In this section, we consider the case when $q$ is odd.
 Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
 $y^{2}=f_{\mathcal{E}}$
-with $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ being a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
+where $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ is a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
 $\sigma(y)=-y.$
@@
 $\det(\varepsilon_{P}(L))=f_{11}^{2}-f_{21}^{2}-(f_{12}^{2}-f_{22}^{2})f_{\mathcal{E}}.$
-For positive integers $k,k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
+For positive integers $k$ and $k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
 $C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{\infty}),f_{2}\in\mathcal{L}(k_{1}Q_{\infty})\}$
@@
 $C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
-where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}\in\mathbb{F}_{q}(x)/\mathbb{F}_{q}$. Through the Singleton bound combined with Theorems IV.1 and IV.2, we derive
+where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}$. By applying the Singleton bound together with Theorems IV.1 and IV.2, we obtain
 $4s-2k\leq 2d_{sr,1},2d_{sr,2}\leq 4s-k+2,$
@@
 Example IV.1.
-Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
+Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ defined over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
 $\{P_{0},P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$
@@
 Since $x^{3}+5$ is irreducible over $\mathbb{F}_{q}[x]$, the function $\det(\varepsilon_{P}(L))$ has no rational zeros. Consequently, the codeword
-$\left(\begin{bmatrix}3,&1\\ -3,&3\end{bmatrix},\ldots,\begin{bmatrix}3,&1\\ 5,&3\end{bmatrix}\right)$
-achieves the maximum sum-rank distance $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
+[…]
+achieves the maximum sum-rank weight of $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
 $\varepsilon_{P}(L)=\begin{bmatrix}x^{3}+1,&0\\ 0,&1-x^{3}\end{bmatrix}$
@@
 All places in $\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$ are zeros of $\det(\varepsilon_{P}(L))$. Thus, the codeword
-$\left(\begin{bmatrix}1,&0\\ 0,&1\end{bmatrix},\ldots,\begin{bmatrix}0,&0\\ 0,&2\end{bmatrix}\right)$
-has sum-rank distance 8, attaining the lower bound.
+[…]
+has a sum-rank weight that attains the lower bound.
 Example IV.2.
@@
 $\{P_{0},P_{2},P_{3},P_{4},P_{5},P_{6}\}.$
-Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
+Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}+P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
 $A_{0}=1,A_{6}=36,A_{7}=144,A_{8}=1542,A_{9}=7944,$
@@
 where $A_{i}$ counts the number of codewords with sum-rank weight $i$.
+IV-D The Gilbert-Varshamov-like Bound
+In this section, we asymptotically compare our codes with the GV-like bound for sum-rank metric codes.
+Suppose that $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}/F_{i}$ is a Galois extension of degree 2 for all $i\in\mathbb{N}$.
+Denote by $\{g_{i}\}_{i\in\mathbb{N}}$ and $\{N(F_{i})\}_{i\in\mathbb{N}}$ the genus and the number of rational places of $F_{i}$, respectively.
+Following the construction from previous sections, with a sequence of integers $\{k_{i}\}_{i\in\mathbb{N}}$ such that $2g_{i}-1\leq k_{i}\geq 2N(F_{i})$, we obtain a sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ with parameters
+  • code length: $s_{i}\leq 4N(F_{i})$,
+  • dimension: $\ell_{k_{i}}=k_{i}-g_{i}+1$,
+  • minimal sum-rank distance: $d_{sr,i}\geq 2s_{i}-k_{i}$.
+Therefore, we have
+$d_{sr,i}+\ell_{k_{i}}\geq 2s_{i}-g_{i}+1.$
+By the Hasse-Weil lower bound over $\mathbb{F}_{q}$, if either $N(F_{i})$ or $g_{i}$ tends to infinity, then so does the other.
+Ihara [20] observed that the Hasse-Weil bound is no longer optimal when the genus of a function field is sufficiently large with respect to $q$. Let $A(q)$ be the Ihara’s constant, defined as
+[…]
+The Drinfeld-Vlǎduţ bound [47] states that $A(q)\leq\sqrt{q}-1$.
+If $q$ is a square, then Drinfeld and Vlǎduţ proved that equality holds.
+Tsfasman, Vlǎduţ, and Zink [45] provided an example of curves that attain this bound employing modular curves.
+Garcia and Stichtenoth [14] presented a more explicit tower of function fields.
+Using their technology, we can present a chain of quadratic function field extensions that attain the bound for some even $q$.
+Proposition IV.1.
+Let $q=2^{2^{h}}$ with $h\geq 2$, and $F_{0}=\mathbb{F}_{q}(z_{0})$ be the rational function field.
+Suppose that $t=2^{h-1}$ and $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}=F_{i}(z_{i+1})$, where $z_{i}$ satisfies the equations:
+  • for $j\geq 0$, $z_{jt+1}^{2}+z_{jt+1}=x_{j}^{2^{t}+1}$, where $x_{0}=z_{0}$, $x_{1}=z_{t}$ and $x_{j}=\frac{z_{jt}}{x_{j-1}}$ for $j\geq 2$,
+  • for $j\geq 0$ and $2\leq i\leq t$, $z_{jt+i}^{2}+z_{jt+i}=z_{jt+i-1}$.
+Then
+$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\sqrt{q}-1.$
+Proof.
+It is sufficient to show that for any $j\geq 0$, the function field $F_{jt}$ has the form $F_{(j+1)t}=F_{jt}(z_{(j+1)t})$, where $z_{(j+1)t}$ satisfies the equation
+$z_{(j+1)t}^{2^{t}}+z_{(j+1)t}=x_{j}^{2^{t}+1},$
+with
+$x_{0}=z_{0},\ x_{1}=z_{t},\ x_{j}=\frac{z_{jt}}{x_{j-1}}\text{ for }j\geq 2.$
+Then, using [14, Corollary 3.2], we obtain
+$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\lim\limits_{j\to\infty}\frac{N_{jt}}{g_{jt}}=\sqrt{q}-1.$
+For $j=0$, observe that
+$z_{t}^{2}+z_{t}=z_{t-1},$
+$z_{t}^{4}+z_{t}^{2}=z_{t-1}^{2},$
+$\cdots$
+$z_{t}^{2^{t}}+z_{t}^{2^{t-1}}=z_{t-1}^{2^{t-1}}.$
+Recall that $t=2^{h-1}$, which implies
+$z_{t}^{2^{t}}+z_{t}$ $=(z_{t-1}^{2^{t-1}}+z_{t-1}^{2^{t-2}})+\cdots+(z_{t-1}^{2}+z_{t-1})$
+$=z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-4}}+\cdots+z_{t-2}^{4}+z_{t-2}$
+$=(z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-3}})+(z_{t-2}^{2^{t-3}}+z_{t-2}^{2^{t-4}})+\cdots+(z_{t-2}^{4}+z_{t-2}^{2})+(z_{t-2}^{2}+z_{t-2})$
+$=z_{t-3}^{2^{t-3}}+z_{t-3}^{2^{t-4}}+\cdots+z_{t-3}^{2}+z_{t-3}$
+$=z_{t-4}^{2^{t-4}}+z_{t-4}^{2^{t-6}}+\cdots+z_{t-4}^{4}+z_{t-4}$
+$=z_{t-5}^{2^{t-5}}+z_{t-5}^{2^{t-6}}+\cdots+z_{t-5}^{2}+z_{t-5}$
+$\qquad\vdots$
+$=z_{1}^{2}+z_{1}$
+$=x_{0}^{2^{t}+1}$
+Then we have $F_{t}=F_{0}(z_{t})$ with
+$z_{t}^{2^{t}}+z_{t}=x_{0}^{2^{t}+1}$
+The proof can be completed with similar equations for any $j\geq 1$. ∎
+The above proposition shows that there exist towers of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound for some square $q$.
+For any sum-rank metric code $C_{i}\in\mathbb{F}_{q}^{2\times 2}$ with $i\in\mathbb{N}$, its rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ are defined as
+$\mathcal{R}_{i}:=\frac{\ell_{k_{i}}}{2s_{i}},\quad\delta_{sr,i}:=\frac{d_{sr,i}}{2s_{i}}.$
+Theorem IV.3.
+Suppose that $q$ is a square and there exists a tower $\{F_{i}\}_{i\in\mathbb{N}}$ of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound.
+For any real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
+$\mathcal{R}_{sr}<1-\delta_{sr}-\frac{1}{2(\sqrt{q}-1)}.$ (1)
+There exists a sum-rank metric code with rate at least $\mathcal{R}_{sr}$ and relative minimum distance at least $\delta_{sr}$.
+Proof.
+We only need to consider the sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ constructed from the tower $\{F_{i}\}_{i\in\mathbb{N}}$ via Theorem IV.1 or IV.2.
+For any $i\in\mathbb{N}$, the parameters of $C_{i}$ are
+$s\leq 4N(F_{i}),\ell_{k_{i}},d_{sr,i}\geq 2s_{i}-k_{i}].$
+Denote the rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ as above.
+Choose $k_{i}$ such that $\mathcal{R}_{i}\geq\mathcal{R}_{sr}$ for all $i$ and $\lim\limits_{i\to\infty}\mathcal{R}_{i}=\mathcal{R}_{sr}$.
+Then we have
+$\mathcal{R}_{sr}=\lim\limits_{i\to\infty}\mathcal{R}_{i}\geq 1-\lim\limits_{i\to\infty}\delta_{sr,i}-\frac{1}{2(\sqrt{q}-1)}.$
+Thus for sufficiently large $i$, the code $C_{i}$ satisfies the desired parameters. ∎
+Note that the GV-like bound for $n=m=2$ has the form
+$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{2}\log_{q}(1+\frac{1}{2\delta_{sr}})-\frac{\log_{q}(1+2\delta_{sr})}{4}$
+$-\frac{\log_{q}(\gamma_{q})}{4}+o(1).$ (2)
+For certain values of $q$ such that there exists a tower of quadratic function field extensions attaining the Drinfeld-Vlǎduţ bound, some families of our sum-rank metric codes surpass the GV-like bound.
+Using the function field tower given in Proposition IV.1, we plot the comparisons between equations (1) with (IV-D) for $q=2^{4}$, $q=2^{8}$ and $q=2^{16}$ in Figs.
+2, 3, and 4, respectively.
+Figure 2: $q=2^{4}$
+Figure 3: $q=2^{8}$
+Figure 4: $q=2^{16}$
 V Conclusion
-In this paper, we conducted two families of $2\times 2$ sum-rank metric codes by exploiting quadratic Kummer extensions of algebraic function fields.
-By representing elements in the extension fields as matrices over the base field through explicit algebraic mappings, we rigorously determined the parameters of the proposed codes, including their minimum sum-rank distances and code rates.
-To demonstrate the practical applicability of our framework, we provided concrete instantiations over elliptic function fields.
-To the best of our knowledge, these code constructions are novel in the context of algebraic-geometric (AG) codes for the sum-rank metric.
-Prior to this work, the only systematic study of AG codes in this metric was established by Berardini and Caruso [4], who proposed a general framework for such codes.
-However, our methodology differs fundamentally from theirs by explicitly exploiting the intrinsic algebraic structures of function field extensions—particularly through matrix representations derived from quadratic Kummer extensions—to control code parameters and optimize distance properties.
-In this paper, we investigate Kummer extensions defined by the equation $y^{2}=u\quad\text{where }u\notin\{\omega^{2}\mid\omega\in F\setminus\mathbb{F}_{q}\}.$ This specific choice of non-square elements $u$ enables us to rigorously establish the non-degeneracy condition $\det(\varepsilon_{P}(L))\neq 0$ in the restricted case where $f_{1}=0$.
-The principal challenge in generalizing these constructions to $n\times n$ matrices stems from the need to preserve the non-vanishing determinant property $\det(\varepsilon_{P}(L))\neq 0\quad\text{for arbitrary }n\geq 2.$ We intend to address this dimensionality constraint through more flexible algebraic frameworks in subsequent work.
-Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension $k-g(F)+1$, we observe the following approximate relationship:
+In this paper, we presented two constructions of $2\times 2$ sum-rank metric codes via quadratic Galois extensions of algebraic function fields.
+By expressing the functions of extension fields to matrices over the base fields, we determined the parameters of these codes.
+Furthermore, we provided illustrative examples based on elliptic function fields.
+To the best of our knowledge, these constructions are new.
+The only existing result on AG codes in the sum-rank metric was given by E.
+Berardini and X.
+Caruso [4].
+In contrast to their work, our approach relies more heavily on the structures of function field extensions.
+In this paper, we established that $\det(\varepsilon_{P}(L))\neq 0$ in the special case where $f_{1}=0$.
+The primary challenge in extending these constructions to $n\times n$ matrices lies in guaranteeing $\det(\varepsilon_{P}(L))\neq 0$.
+Through SageMath computations, we verify that $L(g)$ has trivial kernel for certain special function fields and places when $n\geq 3$.
+Consequently, in our future work, we will investigate more general constructions.
+Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension kg(F)+1k-g(F)+1, we have
 $\max d_{sr}-\min d_{sr}\approx\frac{k+g(F)+1}{2}.$
-This raises an interesting problem: whether techniques developed for constructing MDS AG codes [11] can be adapted to the sum-rank metric setting to achieve MSRD AG codes.
-Following the approach for AG codes with narrow bounds (such as elliptic codes [11]), we propose this problem as an important objective for our subsequent research.
+This raises an interesting problem: adapting the techniques used in constructing MDS AG codes to the sum-rank metric to obtain MSRD AG codes.
+Building on the approach for AG codes with narrow bounds (such as elliptic codes [11]), we identify this problem as an important objective for our subsequent research.
 VI Acknowledgment
-The work of Chang-An Zhao is supported by Guangdong Basic and Applied Basic Research Foundation(No.
-2025A1515011764), Guangdong Major Project of Basic and Applied Basic Research (No.
-2019B030302008), the National Natural Science Foundation of China (No.
+This work is supported by Natural Science Foundation of Guangdong Province (No.
+2025A1515011764), the National Natural Science Foundation of China (No.
 12441107) and Guangdong Provincial Key Laboratory of Information Security Technology (No.